# Grassroots Bonds: A Grassroots Foundation for Market Liquidity

### 14 Mar 2026 v1

@@ Abstract.
Here we introduce grassroots bonds, which extend grassroots coins with a maturity date, reframing grassroots coins—cash—as mature grassroots bonds. Bond redemption generalises coin redemption, allowing the lending of liquid coins in exchange for interest-bearing future-maturity bonds. We show that digital social contracts—voluntary agreements among persons, specified, fulfilled, and enforced digitally—can express the full gamut of financial instruments as the voluntary swap of grassroots bonds, including credit lines, loans, sale of debt, forward contracts, options, and escrow-based instruments, and that classical liquidity ratios are applicable just as well to grassroots bonds. The formal specification presented here was used by AI to derive a working implementation of grassroots bonds in GLP, a concurrent logic programming language implemented in Dart for smartphone deployment. The implementation is illustrated by a running multiagent village market scenario, also implemented in GLP by AI.
@@ Appendix B Bond Agent Source Code
The complete bond agent is approximately 740 lines of typed GLP code. The full source, including type definitions, infrastructure helpers, and all twelve test scenarios, is available at https://github.com/EShapiro2/GLP-Bonds.
%% bond_agent.glp - Bond agent for Grassroots Bonds
%% Extends the social graph agent (agent/4) to agent/6 with:
%%   - Holdings: list of bonds held by this agent
%%   - NextSerial: next serial number for minting
%% Phase 1: mint + balance + done (single-agent, no network)
%% Phase 2: social graph befriending
%% Phase 7: trade (credit, loan, payment,
%%   redemption, sale of debt, etc.)
%% Date: 2026-03-03
@@
%% TYPE DEFINITIONS
%% Stream and Channel are predefined in prelude:
%%   Stream ::= [] ; [_|Stream].
%%   Channel ::= ch(Stream, Stream?).
@@
LotList ::= [] ; [Lot | LotList].
TradeResponse ::= trade_accept(BondList)
               ; trade_decline(BondList)
               ; trade_decline_menu(BondList, BondList).
@@
%% Escrow results -- typed unions for
%%   escrow output variables
EscrowBenResult ::= escrow_bonds(BondList)
@@
EscrowDepResult ::= escrow_bonds(BondList) ; escrow_expired.
@@
                ; trade_propose(LotList,
                    BondList, TradeResponse?)
                ; escrow_offer(Constant,
                    EscrowBenResult).
@@
FriendStream  ::= [] ; [FriendMsg | FriendStream].
FriendChannel ::= ch(FriendStream, FriendStream?).
@@
%% Network content (2-arg cold-call: msg(Dest, Content))
@@
%% --- Content types for messages ---
%% Content carried inside user-to-agent
%%   messages (msg(’_user’, Id, Content))
@@
              ; decision(Decision, Constant,
                  PendingValue)
@@
              ; trade(Constant, LotList, LotList)
@@
              ; deposit_escrow(Constant, LotList, Constant)
@@
               ; trade_pending(TradeResponse?,
                   LotList, BondList)
@@
%% Content carried inside agent-to-user messages
@@
               ; balance_report(BondList)
@@
               ; rejected
               ; rejected(Constant)
               ; trade_proposed(Constant,
                   LotList, TradeResponse?,
                   BondList)
@@
               ; escrow_deposited(Constant,
                   Constant, EscrowCancel?)
@@
               ; trade_returned_menu(Constant, BondList).
%% Agent input stream (from mediator + injected results)
@@
               ; trade_complete(Constant, BondList)
               ; trade_returned_bonds(Constant, BondList)
               ; escrow_ben_released(Constant, BondList)
@@
               ; escrow_dep_returned(Constant, BondList)
               ; trade_returned_bonds_menu(
                   Constant, BondList,
                   BondList).
UserInStream ::= [] ;
    [UserInMsg | UserInStream].
MediatorToAgentMsg ::=
    msg(Constant, Constant, UserContent).
MediatorToAgentStream ::= [] ;
    [MediatorToAgentMsg |
     MediatorToAgentStream].
AgentToUserMsg    ::= msg(Constant, Constant, AgentContent).
AgentToUserStream ::= [] ;
    [AgentToUserMsg | AgentToUserStream].
AgentChannel ::=
    ch(AgentToUserStream,
       MediatorToAgentStream?).
NetInMsg ::= msg(Constant, NetColdCall)
           ; msg(Constant, Constant, FriendContent).
NetInStream ::= [] ; [NetInMsg | NetInStream].
OutputKey   ::= ’_user’ ; ’_net’ ; friend(Constant).
OutputContent ::= minted(Constant, Constant)
                ; balance_report(BondList)
                ; befriend(Constant, Response?)
                ; connected(Constant)
                ; rejected
                ; rejected(Constant)
                ; response(Response)
                ; trade_propose(LotList,
                    BondList, TradeResponse?)
                ; trade_proposed(Constant,
                    LotList, TradeResponse?,
                    BondList)
                ; trade_completed(Constant)
                ; trade_failed(Constant)
                ; trade_returned(Constant)
                ; escrow_offer(Constant,
                    EscrowBenResult)
                ; escrow_deposited(Constant,
                    Constant, EscrowCancel?)
                ; escrow_received(Constant, Constant)
                ; escrow_released(Constant)
                ; escrow_cancelled(Constant)
                ; escrow_expired(Constant)
                ; escrow_returned(Constant)
                ; escrow_failed(Constant)
                ; trade_returned_menu(Constant, BondList).
@@
OutputStream ::= [] ; [OutputMsg | OutputStream].
OutputEntry ::= output(OutputKey, OutputStream?).
OutputsList ::= [] ; [OutputEntry | OutputsList].
%% MERGE_NET_IN
procedure merge_net_in(NetInStream?,
    NetInStream?, NetInStream).
merge_net_in([X|Xs], Ys, [X?|Zs?]) :-
    merge_net_in(Ys?, Xs?, Zs).
merge_net_in(Xs, [Y|Ys], [Y?|Zs?]) :-
    merge_net_in(Xs?, Ys?, Zs).
merge_net_in([], Ys, Ys?).
merge_net_in(Xs, [], Xs?).
%% LOOKUP_SEND
procedure lookup_send(OutputKey?,
    OutputMsg?, OutputsList?,
    OutputsList).
@@
procedure lookup_send_step(OutputKey?,
    OutputMsg?, OutputsList?,
    OutputsList).
lookup_send_step(Key, Msg, [output(K, [Msg?|Out1?])|Rest],
@@
lookup_send_step(Key, Msg, [output(K, Out?)|Rest],
@@
%% ADD_OUTPUT / CLOSE_OUTPUTS
procedure add_output(OutputKey?,
    OutputStream, OutputsList?,
    OutputsList).
@@
procedure close_outputs(OutputsList?).
close_outputs([output(_, [])|Outs]) :- close_outputs(Outs?).
@@
%% INJECT_MSG -- Inject cold-call response into UserIn
@@
    UserInStream?, UserInStream).
@@
%% NEW_FRIEND_CHANNEL
procedure new_friend_channel(FriendChannel, FriendChannel).
new_friend_channel(ch(Xs?, Ys), ch(Ys?, Xs)).
%% RESPONSE HANDLING -- bind cold-call decision
procedure bind_response(Decision?, Constant?, Response,
    OutputsList?, OutputsList, NetInStream?, NetInStream).
bind_response(yes, From, accept(RetCh?),
    Outs, Outs1?, In, In1?) :-
    new_friend_channel(RetCh, LocalCh) |
    handle_response(accept(LocalCh?),
        From?, Outs?, Outs1, In?, In1).
bind_response(no, _, no, Outs, Outs1?, In, In?) :-
    lookup_send(’_user’, msg(agent, ’_user’, rejected),
        Outs?, Outs1).
procedure handle_response(Response?, Constant?,
    OutputsList?, OutputsList, NetInStream?, NetInStream).
handle_response(accept(ch(FIn, FOut?)), From,
    Outs, Outs2?, In, In1?) :-
    ground(From?) |
    add_output(friend(From?), FOut, Outs?, Outs1),
    lookup_send(’_user’,
        msg(agent, ’_user’,
            connected(From?)),
        Outs1?, Outs2),
    merge_net_in(In?, FIn?, In1).
handle_response(no, From, Outs, Outs1?, In, In?) :-
    ground(From?) |
    lookup_send(’_user’,
        msg(agent, ’_user’,
            rejected(From?)),
        Outs?, Outs1).
%% CREATE_BONDS
procedure create_bonds(Constant?,
    Constant?, Constant?, Constant?,
    BondList).
@@
%% APPEND
procedure append(BondList?, BondList?, BondList).
append([X|Xs], Ys, [X?|Zs?]) :- append(Xs?, Ys?, Zs).
@@
%% SELECT_BONDS_EXACT
procedure select_bonds_exact(Constant?,
    Constant?, Constant?,
    BondList?, Constant,
    BondList, BondList).
@@
    [bond(I, M, S)|Rest],
    Status?, [bond(I?, M?, S?)|Sel?], Rem?) :-
@@
    select_bonds_exact(Issuer?, Maturity?, K1?, Rest?,
        Status, Sel, Rem).
@@
    select_bonds_exact(Issuer?, Maturity?, K?, Rest?,
        Status, Sel, Rem).
@@
%% SELECT_BONDS_BY_SPEC -- Fold over LotList
procedure select_bonds_by_spec(LotList?,
    BondList?, Constant,
    BondList, BondList).
@@
    [lot(Issuer, Maturity, Count)|Rest],
    Holdings, Status?, AllSel?,
    FinalRem?) :-
@@
    select_bonds_exact(Issuer?, Maturity?,
        Count?, Holdings?,
        LotStatus, LotSel, LotRem),
    select_by_spec_continue(LotStatus?,
        Rest?, LotSel?, LotRem?,
@@
select_bonds_by_spec([], Holdings, ok, [], Holdings?).
procedure select_by_spec_continue(
    Constant?, LotList?, BondList?,
    BondList?, Constant,
    BondList, BondList).
@@
    select_bonds_by_spec(Rest?, LotRem?, Status, RestSel,
        FinalRem),
@@
%% BIND_TRADE -- Bind TradeResponse variable (writer mode)
procedure bind_trade_accept(TradeResponse, BondList?).
@@
procedure bind_trade_decline(TradeResponse, BondList?).
@@
%% INJECT_TRADE_RESULT
procedure inject_trade_result(TradeResponse?, Constant?,
    UserInStream?, UserInStream).
inject_trade_result(trade_accept(TheirBonds), From, Ys,
    [trade_complete(From?, TheirBonds?)|Ys?]) :-
@@
inject_trade_result(trade_decline(OurBonds), From, Ys,
@@
    trade_decline_menu(OurBonds, Menu),
    From, Ys,
    [trade_returned_bonds_menu(
        From?, OurBonds?, Menu?)|Ys?]) :-
@@
inject_trade_result(Resp, From, [Y|Ys], [Y?|Ys1?]) :-
@@
%% DO_TRADE -- Select bonds matching
%%   GiveSpec, then send or fail
procedure do_trade(Constant?, Constant?, LotList?, LotList?,
    BondList?, UserInStream?, NetInStream?, OutputsList?,
    Constant?).
do_trade(Id, Target, GiveSpec, WantSpec, Holdings,
    UserIn, NetIn, Outs, NextSerial) :-
    select_bonds_by_spec(GiveSpec?, Holdings?,
        Status, Selected, Remaining),
    do_trade_result(Status?, Id?, Target?, WantSpec?,
        Selected?, Remaining?,
        UserIn?, NetIn?, Outs?, NextSerial?).
procedure do_trade_result(Constant?, Constant?, Constant?,
    LotList?, BondList?, BondList?, UserInStream?,
    NetInStream?, OutputsList?, Constant?).
do_trade_result(ok, Id, Target, WantSpec,
    Selected, Remaining,
    UserIn, NetIn, Outs, NextSerial) :-
    lookup_send(friend(Target?),
        msg(Id?, Target?,
            trade_propose(WantSpec?,
                Selected?, TradeResp)),
        Outs?, Outs1),
    inject_trade_result(TradeResp?,
        Target?, UserIn?, UserIn1),
    agent(Id?, UserIn1?, NetIn?, Outs1?, Remaining?,
        NextSerial?).
do_trade_result(fail, Id, Target, _, Selected, Remaining,
    UserIn, NetIn, Outs, NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(Target?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?, OrigHoldings?,
        NextSerial?).
%% HANDLE_TRADE_FILL
procedure handle_trade_fill(Constant?, Constant?, Constant?,
    TradeResponse, BondList?, BondList?, BondList?,
    UserInStream?, NetInStream?, OutputsList?, Constant?).
handle_trade_fill(ok, Id, From, trade_accept(Selected?),
    OfferedBonds, Selected, Remaining,
    UserIn, NetIn, Outs, NextSerial) :-
    append(Remaining?, OfferedBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_completed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
handle_trade_fill(fail, Id, From,
    trade_decline(OfferedBonds?), OfferedBonds,
    Selected, Remaining,
    UserIn, NetIn, Outs, NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?, OrigHoldings?,
        NextSerial?).
%% CLASSIFY_TRADE -- Classify incoming trade for auto-accept
procedure classify_trade(Constant?,
    LotList?, BondList?, Constant).
@@
classify_trade(_, _, _, normal) :- otherwise | true.
%% BUILD_MENU -- One bond per issuer, earliest maturity
procedure build_menu(Constant?, BondList?, BondList).
@@
procedure build_menu_acc(Constant?,
    BondList?, BondList?, BondList).
build_menu_acc(Id, [bond(I, _, _)|Rest], Acc, Menu?) :-
@@
build_menu_acc(Id, [bond(I, M, S)|Rest], Acc, Menu?) :-
@@
procedure menu_update(Constant?, Constant?, Constant?,
    BondList?, BondList).
@@
%% TRADE_DISPATCH
procedure trade_dispatch(Constant?, Constant?, Constant?,
    LotList?, BondList?, TradeResponse,
    BondList?, UserInStream?, NetInStream?,
    OutputsList?, Constant?).
@@
    trade_accept([]),
    Holdings, UserIn, NetIn, Outs, NextSerial) :-
@@
    agent(Id?, UserIn?, NetIn?, Outs1?, NewHoldings?,
        NextSerial?).
trade_dispatch(redemption, Id, From,
    WantSpec, OfferedBonds, TradeResp?,
    Holdings, UserIn, NetIn, Outs, NextSerial) :-

### 17 Mar 2026 v2

@@ Abstract.
Here we introduce grassroots bonds, which extend grassroots coins with a maturity date, reframing grassroots coins—cash—as mature grassroots bonds. Bond redemption generalises coin redemption, allowing the lending of liquid coins in exchange for interest-bearing future-maturity bonds. We show that digital social contracts—voluntary agreements among persons, specified, fulfilled, and enforced digitally—can express the full gamut of financial instruments as the voluntary swap of grassroots bonds, including credit lines, loans, sale of debt, forward contracts, options, and escrow-based instruments, and that classical liquidity ratios are applicable just as well to grassroots bonds. Grassroots bonds may thus allow local digital economies to form and grow without initial capital or external credit, harnessing mutual trust within communities into liquidity.
The formal specification presented here was used by AI to derive a working implementation of grassroots bonds in GLP, a concurrent logic programming language implemented in Dart for smartphone deployment. The implementation is illustrated by a running multiagent village market scenario, also implemented in GLP by AI.
@@ Appendix B Bond Agent Source Code
The bond agent implementation consists of two GLP modules: self.glp, which defines all shared types and helper procedures visible to all modules via ancestor scoping; and agent.glp, which defines the bond agent and its local procedures. Together they comprise approximately 740 lines of typed GLP code. The full source, including the UI mediator, twelve test scenarios, and the six village market actors, is available at https://github.com/EShapiro2/GLP-Bonds.
Module 1: Shared Types and Helpers (self.glp)
%% self.glp — Bonds V2 shared type vocabulary
%% All protocol types for the Grassroots Bonds application.
%% Defined once here, visible to all modules via ancestor scoping.
-module(bonds).
@@
%% Stream(X) and Channel(In, Out) are predefined in prelude:
%%   Stream(X) ::= [] ; [X | Stream(X)].
%%   Channel(In, Out) ::= ch(In, Out?).
@@
TradeResponse ::= trade_accept(Stream(Bond))
               ; trade_decline(Stream(Bond))
               ; trade_decline_menu(Stream(Bond), Stream(Bond)).
@@
%% Escrow results
EscrowBenResult ::= escrow_bonds(Stream(Bond))
@@
EscrowDepResult ::= escrow_bonds(Stream(Bond))
                  ; escrow_expired.
@@
                ; trade_propose(Stream(Lot), Stream(Bond),
                    TradeResponse?)
                ; escrow_offer(Constant, EscrowBenResult).
@@
FriendChannel ::= ch(Stream(FriendMsg),
                     Stream(FriendMsg)?).
@@
%% Network content (2-arg cold-call)
@@
%% User-to-agent message content
@@
  ; decision(Decision, Constant, PendingValue)
@@
  ; trade(Constant, Stream(Lot), Stream(Lot))
@@
  ; deposit_escrow(Constant, Stream(Lot), Constant)
@@
  ; trade_pending(TradeResponse?, Stream(Lot),
      Stream(Bond))
@@
%% Agent-to-user message content
@@
  ; balance_report(Stream(Bond))
@@
  ; rejected ; rejected(Constant)
  ; trade_proposed(Constant, Stream(Lot),
      TradeResponse?, Stream(Bond))
@@
  ; escrow_deposited(Constant, Constant,
      EscrowCancel?)
@@
  ; trade_returned_menu(Constant, Stream(Bond)).
AgentToUserMsg     ::= msg(Constant, Constant,
                           AgentContent).
MediatorToAgentMsg ::= msg(Constant, Constant,
                           UserContent).
%% Network input message
NetInMsg ::= msg(Constant, NetColdCall)
           ; msg(Constant, Constant, FriendContent).
%% Agent input stream (mediator + injected results)
@@
  ; trade_complete(Constant, Stream(Bond))
  ; trade_returned_bonds(Constant, Stream(Bond))
  ; escrow_ben_released(Constant, Stream(Bond))
@@
  ; escrow_dep_returned(Constant, Stream(Bond))
  ; trade_returned_bonds_menu(Constant,
      Stream(Bond), Stream(Bond)).
%% Output types (type union of agent + friend content)
OutputContent ::= AgentContent ; FriendContent.
@@
OutputKey   ::= ’_user’ ; ’_net’ ; friend(Constant).
OutputEntry ::= output(OutputKey, Stream(OutputMsg)?).
AgentChannel ::= ch(Stream(AgentToUserMsg),
                    Stream(MediatorToAgentMsg)?).
%% Shared helper procedures
procedure merge(Stream(X)?, Stream(X)?, Stream(X)).
merge([X|Xs], Ys, [X?|Zs?]) :-
    merge(Ys?, Xs?, Zs).
merge(Xs, [Y|Ys], [Y?|Zs?]) :-
    merge(Xs?, Ys?, Zs).
merge([], Ys, Ys?).
merge(Xs, [], Xs?).
procedure lookup_send(OutputKey?, OutputMsg?,
    Stream(OutputEntry)?, Stream(OutputEntry)).
@@
procedure lookup_send_step(OutputKey?, OutputMsg?,
    Stream(OutputEntry)?, Stream(OutputEntry)).
lookup_send_step(Key, Msg,
    [output(K, [Msg?|Out1?])|Rest],
@@
lookup_send_step(Key, Msg,
    [output(K, Out?)|Rest],
@@
procedure add_output(OutputKey?, Stream(OutputMsg),
    Stream(OutputEntry)?, Stream(OutputEntry)).
@@
procedure close_outputs(Stream(OutputEntry)?).
close_outputs([output(_, [])|Outs]) :-
    close_outputs(Outs?).
@@
    Stream(UserInMsg)?, Stream(UserInMsg)).
@@
procedure create_bonds(Constant?, Constant?,
    Constant?, Constant?, Stream(Bond)).
@@
procedure append(Stream(Bond)?, Stream(Bond)?,
    Stream(Bond)).
append([X|Xs], Ys, [X?|Zs?]) :-
    append(Xs?, Ys?, Zs).
@@
procedure select_bonds_exact(Constant?, Constant?,
    Constant?, Stream(Bond)?, Constant,
    Stream(Bond), Stream(Bond)).
@@
    [bond(I, M, S)|Rest], Status?,
    [bond(I?, M?, S?)|Sel?], Rem?) :-
@@
    select_bonds_exact(Issuer?, Maturity?, K1?,
        Rest?, Status, Sel, Rem).
@@
    select_bonds_exact(Issuer?, Maturity?, K?,
        Rest?, Status, Sel, Rem).
@@
procedure select_bonds_by_spec(Stream(Lot)?,
    Stream(Bond)?, Constant,
    Stream(Bond), Stream(Bond)).
@@
    [lot(Issuer, Maturity, Count)|Rest], Holdings,
    Status?, AllSel?, FinalRem?) :-
@@
    select_bonds_exact(Issuer?, Maturity?, Count?,
        Holdings?, LotStatus, LotSel, LotRem),
    select_by_spec_continue(LotStatus?, Rest?,
        LotSel?, LotRem?,
@@
select_bonds_by_spec([], Holdings, ok, [],
    Holdings?).
procedure select_by_spec_continue(Constant?,
    Stream(Lot)?, Stream(Bond)?, Stream(Bond)?,
    Constant, Stream(Bond), Stream(Bond)).
@@
    select_bonds_by_spec(Rest?, LotRem?,
        Status, RestSel, FinalRem),
@@
procedure bind_trade_accept(TradeResponse,
    Stream(Bond)?).
@@
procedure bind_trade_decline(TradeResponse,
    Stream(Bond)?).
@@
procedure inject_trade_result(TradeResponse?,
    Constant?, Stream(UserInMsg)?,
    Stream(UserInMsg)).
inject_trade_result(trade_accept(TheirBonds), From,
    Ys, [trade_complete(From?, TheirBonds?)|Ys?]) :-
@@
inject_trade_result(trade_decline(OurBonds), From,
    Ys,
@@
    trade_decline_menu(OurBonds, Menu), From, Ys,
    [trade_returned_bonds_menu(From?, OurBonds?,
        Menu?)|Ys?]) :-
@@
inject_trade_result(Resp, From, [Y|Ys],
    [Y?|Ys1?]) :-
@@
procedure classify_trade(Constant?, Stream(Lot)?,
    Stream(Bond)?, Constant).
@@
classify_trade(_, _, _, normal) :-
    otherwise | true.
procedure build_menu(Constant?, Stream(Bond)?,
    Stream(Bond)).
@@
procedure build_menu_acc(Constant?, Stream(Bond)?,
    Stream(Bond)?, Stream(Bond)).
build_menu_acc(Id, [bond(I, _, _)|Rest], Acc,
    Menu?) :-
@@
build_menu_acc(Id, [bond(I, M, S)|Rest], Acc,
    Menu?) :-
@@
procedure menu_update(Constant?, Constant?,
    Constant?, Stream(Bond)?, Stream(Bond)).
@@
procedure escrow(Constant?, Stream(Bond)?,
    EscrowCancel?, EscrowBenResult, EscrowDepResult).
escrow(T, Bonds, _,
    escrow_bonds(Bonds?), escrow_expired) :-
    wait_until(T?) | true.
escrow(_, Bonds, cancel,
    escrow_cancelled, escrow_bonds(Bonds?)).
procedure inject_escrow_ben_result(EscrowBenResult?,
    Constant?, Stream(UserInMsg)?,
    Stream(UserInMsg)).
inject_escrow_ben_result(escrow_bonds(Bonds), From,
    Ys,
    [escrow_ben_released(From?, Bonds?)|Ys?]) :-
    ground(From?), ground(Bonds?) | true.
inject_escrow_ben_result(escrow_cancelled, From, Ys,
    [escrow_ben_cancelled(From?)|Ys?]) :-
    ground(From?) | true.
inject_escrow_ben_result(Resp, From, [Y|Ys],
    [Y?|Ys1?]) :-
    inject_escrow_ben_result(Resp?, From?,
        Ys?, Ys1).
procedure inject_escrow_dep_result(EscrowDepResult?,
    Constant?, Stream(UserInMsg)?,
    Stream(UserInMsg)).
inject_escrow_dep_result(escrow_bonds(Bonds), Target,
    Ys,
    [escrow_dep_returned(Target?, Bonds?)|Ys?]) :-
    ground(Target?), ground(Bonds?) | true.
inject_escrow_dep_result(escrow_expired, Target, Ys,
    [escrow_dep_expired(Target?)|Ys?]) :-
    ground(Target?) | true.
inject_escrow_dep_result(Resp, Target, [Y|Ys],
    [Y?|Ys1?]) :-
    inject_escrow_dep_result(Resp?, Target?,
        Ys?, Ys1).
procedure bind_escrow_cancel(EscrowCancel).
bind_escrow_cancel(cancel).
procedure lookup_pending(ReqId?, PendingValue,
    Stream(PendingEntry)?, Stream(PendingEntry)).
lookup_pending(ReqId, Val?,
    [pending(Id, Val) | Rest], Rest?) :-
    ReqId? =?= Id? | true.
lookup_pending(ReqId, Val?, [P | Rest],
    [P? | Rest1?]) :-
    otherwise |
    lookup_pending(ReqId?, Val, Rest?, Rest1).
lookup_pending(_, error, [], []).
Module 2: Bond Agent (agent.glp)
%% agent.glp — Bond agent module (v2)
%% Exports: agent/6
%% Types and shared helpers: from self.glp
-module(agent).
-mode(system).
%% Merge — local copy (Stream(X) subtyping limitation)
procedure merge(Stream(X)?, Stream(X)?, Stream(X)).
merge([X|Xs], Ys, [X?|Zs?]) :-
    merge(Ys?, Xs?, Zs).
merge(Xs, [Y|Ys], [Y?|Zs?]) :-
    merge(Xs?, Ys?, Zs).
merge([], Ys, Ys?).
merge(Xs, [], Xs?).
%% Response handling — local (calls merge)
procedure bind_response(Decision?, Constant?,
    Response, Stream(OutputEntry)?,
    Stream(OutputEntry), Stream(NetInMsg)?,
    Stream(NetInMsg)).
bind_response(yes, From, accept(RetCh?),
    Outs, Outs1?, In, In1?) :-
    new_channel(RetCh, LocalCh) |
    handle_response(accept(LocalCh?), From?,
        Outs?, Outs1, In?, In1).
bind_response(no, _, no, Outs, Outs1?, In, In?) :-
    lookup_send(’_user’,
        msg(agent, ’_user’, rejected), Outs?, Outs1).
procedure handle_response(Response?, Constant?,
    Stream(OutputEntry)?, Stream(OutputEntry),
    Stream(NetInMsg)?, Stream(NetInMsg)).
handle_response(accept(ch(FIn, FOut?)), From,
    Outs, Outs2?, In, In1?) :-
    ground(From?) |
    add_output(friend(From?), FOut, Outs?, Outs1),
    lookup_send(’_user’,
        msg(agent, ’_user’, connected(From?)),
        Outs1?, Outs2),
    merge(In?, FIn?, In1).
handle_response(no, From, Outs, Outs1?, In, In?) :-
    ground(From?) |
    lookup_send(’_user’,
        msg(agent, ’_user’, rejected(From?)),
        Outs?, Outs1).
%% Trade — select bonds then send or fail
procedure do_trade(Constant?, Constant?,
    Stream(Lot)?, Stream(Lot)?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
do_trade(Id, Target, GiveSpec, WantSpec, Holdings,
    UserIn, NetIn, Outs, NextSerial) :-
    select_bonds_by_spec(GiveSpec?, Holdings?,
        Status, Selected, Remaining),
    do_trade_result(Status?, Id?, Target?,
        WantSpec?, Selected?, Remaining?,
        UserIn?, NetIn?, Outs?, NextSerial?).
procedure do_trade_result(Constant?, Constant?,
    Constant?, Stream(Lot)?, Stream(Bond)?,
    Stream(Bond)?, Stream(UserInMsg)?,
    Stream(NetInMsg)?, Stream(OutputEntry)?,
    Constant?).
do_trade_result(ok, Id, Target, WantSpec, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    lookup_send(friend(Target?),
        msg(Id?, Target?,
            trade_propose(WantSpec?, Selected?,
                TradeResp)),
        Outs?, Outs1),
    inject_trade_result(TradeResp?, Target?,
        UserIn?, UserIn1),
    agent(Id?, UserIn1?, NetIn?, Outs1?,
        Remaining?, NextSerial?).
do_trade_result(fail, Id, Target, _, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(Target?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Fill trade or decline
procedure handle_trade_fill(Constant?, Constant?,
    Constant?, TradeResponse, Stream(Bond)?,
    Stream(Bond)?, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
handle_trade_fill(ok, Id, From,
    trade_accept(Selected?), OfferedBonds, Selected,
    Remaining, UserIn, NetIn, Outs, NextSerial) :-
    append(Remaining?, OfferedBonds?, NewHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_completed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
handle_trade_fill(fail, Id, From,
    trade_decline(OfferedBonds?), OfferedBonds,
    Selected, Remaining, UserIn, NetIn, Outs,
    NextSerial) :-
    append(Selected?, Remaining?, OrigHoldings),
    lookup_send(’_user’,
        msg(agent, ’_user’, trade_failed(From?)),
        Outs?, Outs1),
    agent(Id?, UserIn?, NetIn?, Outs1?,
        OrigHoldings?, NextSerial?).
%% Trade dispatch — auto-accept vs. user decision
procedure trade_dispatch(Constant?, Constant?,
    Constant?, Stream(Lot)?, Stream(Bond)?,
    TradeResponse, Stream(Bond)?,
    Stream(UserInMsg)?, Stream(NetInMsg)?,
    Stream(OutputEntry)?, Constant?).
@@
    trade_accept([]), Holdings, UserIn, NetIn,
    Outs, NextSerial) :-
@@
    agent(Id?, UserIn?, NetIn?, Outs1?,
        NewHoldings?, NextSerial?).
trade_dispatch(redemption, Id, From, WantSpec,
    OfferedBonds, TradeResp?, Holdings, UserIn,
    NetIn, Outs, NextSerial) :-